I am asking for your help with this problem because I can't even begin to solve it.Thanks for your time and what you can help me out with. I am looking for a math tutor that can actually spend time with me for my summer course.If anyone is available please let me know.Thanks
Assume f(x) = x2 + 2. Find the functional value requested.
f(-2) =

Answer by jorel1380(3719)

f(-2)=(-2)2+2=4+2=6..